40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 15 10:52:38 2019, gautam, Update, SUS, EY door opened 
    Reply  Tue Jan 15 15:49:47 2019, gautam, Update, SUS, EY door opened EY_OSEMs.png
       Reply  Sat Jan 19 15:33:18 2019, gautam, Update, SUS, ETMY OSEMs faulty Screen_Shot_2019-01-19_at_3.32.35_PM.pngETMY_sensors_1231832635.pdf
          Reply  Tue Jan 22 20:36:53 2019, gautam, Update, SUS, ETMY OSEMs faulty 
             Reply  Wed Jan 23 12:39:18 2019, gautam, Update, SUS, EY chamber work 6x
                Reply  Wed Jan 23 23:12:44 2019, gautam, Update, SUS, Prep for FC cleaning IY_level_before.pdf
                   Reply  Thu Jan 24 15:32:31 2019, gautam, Update, SUS, Y arm cavity side first contact applied Yarm_FC.pdfOSEMfilter.png
                      Reply  Tue Jan 29 22:12:40 2019, gautam, Update, SUS, Alignment prep 
                         Reply  Wed Jan 30 11:54:24 2019, gautam, Update, SUS, More alignment prep IMG_5958.JPGIMG_5962.JPG
                            Reply  Wed Jan 30 19:25:40 2019, gautam, Update, SUS, Xarm cavity alignment Xarm.pngXarm_IR.png
                               Reply  Fri Feb 1 01:24:06 2019, gautam, Update, SUS, Almost ready for pumpdown tomorrow 
                                  Reply  Fri Feb 1 13:16:50 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                     Reply  Fri Feb 1 21:52:57 2019, gautam, Update, SUS, Pumpdown 83 underway 
                                  Reply  Fri Feb 1 14:44:14 2019, gautam, Update, SUS, Y arm FC cleaning and reinstall IMG_5974.JPGIMG_5986.JPGIMG_5992.JPG
    Reply  Wed Jan 16 16:25:25 2019, gautam, Update, SUS, Yarm locked Yarm_locked.png
       Reply  Fri Jan 18 21:34:18 2019, gautam, Update, SUS, Unused optic on EY table IMG_5408.JPG
          Reply  Sat Jan 19 05:07:45 2019, Koji, Update, SUS, Unused optic on EY table 
Message ID: 14413     Entry time: Wed Jan 23 12:39:18 2019     In reply to: 14411     Reply to this: 14415
Author: gautam 
Type: Update 
Category: SUS 
Subject: EY chamber work 

While Chub is making new cables for the EY satellite box...

  1. I removed the unused optic on the NW corner of the EY table. It is stored in a clean Al-foil lined plastic box, and will be moved to the clean hardware section of the lab (along the South arm, south of MC2 chamber).
  2. Checked table leveling - Attachment #1, looked good, and has been stable over the weekend.
  3. I moved the two oversized washers on the reflector, which I believe are only used because the screw is long and wouldn't go in all the way otherwise. As shown in Attachment #2, this reduces the risk of clipping the main IFO beam axis.
  4. Yesterday, I pulled up the 40m CAD drawing, and played around with a rectangular box that approximates the extents of the elliptical reflector, to see what would be a good place to put it. I chose to go ahead with Attachment #3. Also shown is the eventually realized layout. Note that we'd actually like the dimension marked ~7.6 inches to be more like 7.1 inches, so the optic is actually ~0.5 inch ahead of the second focus of the ellipse, but I think this is good enough. 
  5. Attachment #4 shows the view of the optic as seen from the aperture on the back of the elliptical reflector. Looks good to me.
  6. Having positioned the reflector, I then inserted the heater into the aperture such that it is ~2/3rds the way in, which was the best position found by Annalisa last summer. I then ran 0.9 A of current through the heater for ~ 5 minutes. Attachment #5 shows the optic as seen with the FLIR with no heating, and after 5 minutes of heating. I'd say this is pretty unambiguous evidence that we are indeed heating the mirror. The gradient shown is significantly less pronounced than in Annalisa's simulations (~3K as opposed to 10K), but maybe the FLIR calibration isn't so great.
  7. For completeness, Attachment #6 shows the leveling of the table after this work. Nothing has chanegd significantly.

While the position of the reflector could possibly be optimized further, since we are already seeing a temperature gradient on the optic, I propose pushing on with other vent activities. I'm almost certain the current positioning places the optic closer to the second focus, and we already saw shifts of the HOM resonances with the old configuration, so I'd say we run with this and revisit if needed.

If Chub gives the Sat. Box the green flag, we will work on F.C.ing the mirrors in the evening, with the aim of closing up tomorrow/Friday. 

All raw images in this elog have been uploaded to the 40m google photos.

Attachment 1: leveling.pdf  13.119 MB  | Hide | Hide all | Show all
leveling.pdf
Attachment 2: IMG_5930.jpg  6.295 MB  | Hide | Hide all | Show all
IMG_5930.jpg
Attachment 3: Ellipse_layout.pdf  5.397 MB  | Hide | Hide all | Show all
Ellipse_layout.pdf
Attachment 4: IMG_5932.jpg  5.987 MB  | Hide | Hide all | Show all
IMG_5932.jpg
Attachment 5: hotMirror.pdf  2.195 MB  Uploaded Wed Jan 23 14:12:11 2019  | Show | Hide all | Show all
Attachment 6: EY_leveling_after.pdf  12.363 MB  Uploaded Wed Jan 23 23:12:19 2019  | Hide | Hide all | Show all
EY_leveling_after.pdf
ELOG V3.1.3-